OpenStudy (anonymous):

Using the variation of parameters (Diffeq): y''+4y'+4y=t^(-2)*e^(-2t) I've found the roots, the Wronskian , and set up for my particular solution leaving me with: -e^(-2t)*ln(abs(t))+e^(-2t) However the answer key only has -e^(-2t)*ln(abs(t)) I cant figure out how to get rid of that last term

OpenStudy (anonymous):

\[y''+4y'+4y=0\] gives the characteristic equation \[r^2+4r+4=(r+2)^2=0\] which gives homogeneous solutions \(y_1=C_1e^{-2t}\) and \(C_2te^{-2t}\). Checking for linear independence: \[\begin{align*}W(y_1,y_2)&=\begin{vmatrix}e^{-2t}&te^{-2t}\\-2e^{-2t}&(1-2t)e^{-2t}\end{vmatrix}\\\\ &=(1-2t)e^{-4t}+2te^{-4t}\\\\ &=e^{-4t}\\\\ &\not=0 \end{align*}\] Factors given by VoP: \[\begin{align*}u_1&=-\int\frac{y_2t^{-2}e^{-2t}}{W(y_1,y_2)}~dt\\\\ &=-\int\frac{dt}{t}\\\\ &=-\ln|t| \end{align*}\] \[\begin{align*}u_2&=\int\frac{y_1t^{-2}e^{-2t}}{W(y_1,y_2)}~dt\\\\ &=\int\frac{dt}{t^2}\\\\ &=-\frac{1}{t} \end{align*}\] which means the nonhomogeneous solution is \[\begin{align*}y_p&=u_1y_1+u_2y_2\\\\ &=-e^{-2t}\ln|t|-\frac{e^{-2t}}{t} \end{align*}\]

OpenStudy (anonymous):

Oops, hold on, the final solution is indeed the same as yours. I meant to write \[y_p=-e^{-2t}\ln|t|-e^{-2t}\] The reason it disappears is because of the homogeneous solution, \(y_1=Ce^{-2t}\). The solution \(u_2y_2=e^{-2t}\) gets absorbed into \(y_1\).

OpenStudy (anonymous):

In other words, \[Ce^{-2t}+e^{-2t}=(C+1)e^{-2t}\implies Ce^{-2t}\]

OpenStudy (anonymous):

Basically, I can assume that because \[e ^{-2t}\] is present in all parts, I can just assign a constant to that and because it is the general solution just disregard it? I remember my professor mentioning that now that I see it again, but quite honestly both his accent and handwriting are to hard to understand to get much out of it.